# Self-Data Distillation for Recovering Quality in Pruned Large Language Models

### v2

@@ Abstract
Large language models have driven significant progress in natural language processing, but their deployment requires substantial compute and memory resources. As models scale, compression techniques become essential for balancing model quality with computational efficiency. Structured pruning, which removes less critical components of the model, is a promising strategy for reducing complexity. However, one-shot pruning often results in significant quality degradation, particularly in tasks requiring multi-step reasoning. To recover lost quality, supervised fine-tuning (SFT) is commonly applied, but it can lead to catastrophic forgetting by shifting the model’s learned data distribution. Therefore, addressing the degradation from both pruning and SFT is essential to preserve the original model’s quality. In this work, we propose self-data distilled fine-tuning to address these challenges. Our approach leverages the original, unpruned model to generate a distilled dataset that preserves semantic richness and mitigates catastrophic forgetting by maintaining alignment with the base model’s knowledge. Empirically, we demonstrate that self-data distillation consistently outperforms standard SFT, improving average accuracy by up to 8% on the HuggingFace OpenLLM Leaderboard v1. Specifically, when pruning 6 decoder blocks on Llama3.1-8B Instruct (i.e., 32 to 26 layers, reducing the model size from 8.03B to 6.72B parameters), our method retains 91.2% of the original model’s accuracy compared to 81.7% with SFT, while reducing real-world FLOPs by 16.30%. Furthermore, our approach scales effectively across datasets, with the quality improving as the dataset size increases.
@@ 1 Introduction
The advent of large language models (LLMs) such as GPT-4 (OpenAI et al., 2024), Gemini (Gemini et al., 2024), and Llama 3 (Dubey et al., 2024) has revolutionized natural language processing (NLP), driving significant advancements across various tasks through extensive pre-training on textual data. These models, enhanced by supervised fine-tuning (SFT), demonstrate impressive instruction-following abilities (Ouyang et al., 2022; Touvron et al., 2023a), but come with high computational costs for both training and inference (Kaplan et al., 2020; Hoffmann et al., 2022). To address diverse deployment requirements across varying model scales, sizes, and compute budgets, compressing models for efficient inference is essential, particularly given the significant time, data, and resource constraints associated with training multiple multi-billion parameter models from scratch.
Most model compression techniques can be grouped into four main categories: distillation (Hinton et al., 2015), low-rank factorization (Hu et al., 2022), pruning (LeCun et al., 1989), and quantization (Han et al., 2015). In our work, we focus on pruning, though we aim for our methodology to inspire further developments across these other compression methods. Structured pruning, which selectively removes less critical components of a neural network, has emerged as a promising method for improving LLM efficiency (Ma et al., 2023). This method has gained attention for its ability to reduce memory and computation requirements, making inference more efficient. Recent studies have shown that LLMs exhibit significant redundancy, particularly in the middle layers, where removing these layers has a minimal impact on overall model quality (Men et al., 2024; Gromov et al., 2024). The Transformer (Vaswani et al., 2017) architecture’s residual connections allow for the final output to be a summation of earlier layers, enabling the removal of non-essential layers without drastically harming model quality.
@@
Despite its potential advantages, depth-wise structured pruning presents inherent challenges. It often leads to accuracy degradation, especially on tasks requiring multi-step reasoning, such as ARC-C (Clark et al., 2018) or GSM8k (Cobbe et al., 2021), where the structured order of layer outputs plays a crucial role. In these cases, pruning disrupts the flow of information between layers, resulting in poor model quality even after supervised fine-tuning (SFT) (Sun et al., 2024). While SFT can help recover some of the lost quality, it is generally insufficient for tasks with high reasoning complexity, where the structured sequence of layer outputs is essential. In addition, fine-tuning can amplify catastrophic forgetting (McCloskey and Cohen, 1989), where the model loses previously learned information, particularly on tasks not represented in the fine-tuning data. Standard mitigation strategies, such as data replay (Ostapenko et al., 2022) or parameter importance-based methods (Kirkpatrick et al., 2017), often become impractical for LLMs due to their scale. Moreover, fine-tuning often leads to distribution shifts, further degrading model quality (Yang et al., 2024). As LLMs continue to grow in size and complexity, developing more effective strategies to mitigate these challenges during pruning is critical to unlocking its full potential.
@@
In our work, we propose a novel approach to mitigate the adverse effects of structured pruning by employing self-data distilled fine-tuning. Our method leverages the original, unpruned model as a seed language model to generate a distilled dataset that upholds semantic equivalence with the original task dataset. This approach not only preserves the semantic richness of the data but also mitigates catastrophic forgetting, a phenomenon where fine-tuned models lose their general instruction-following abilities due to the distribution shift introduced during standard SFT. As seen in Figure 1, we show that self-data distillation improves average accuracy by up to 8% over SFT on the HuggingFace OpenLLM Leaderboard v1 (Beeching et al., 2023). Specifically, when pruning 6 blocks from Llama3.1-8B Instruct, our approach retains 91.2% of the original model’s accuracy compared to 81.7% with SFT, while also reducing FLOPs by 16.30%. The main contributions of our work are:
@@
  • To our knowledge, we are the first to introduce self-data distillation as a fine-tuning method for recovering the model quality of pruned models. Empirically, we show that self-data distillation on Llama3.1-8B Instruct consistently outperforms SFT across all pruned models.
@@ 2.1 Layer-Pruning Algorithm for Language Models
Transformer networks (Vaswani et al., 2017) have become a foundational architecture in deep learning, particularly for tasks involving natural language processing and sequence modeling. A standard Transformer consists of $L$ layers, each of which includes a multi-head self-attention mechanism followed by a feedforward network. These layers sequentially transform an input sequence into increasingly abstract representations, with each layer’s output serving as the input to the subsequent layer. Let $x^{(\ell)}$ denote the input to the $\ell^{th}$ layer, and $h^{(\ell)}=f(x^{(\ell)})$ denote the output of the $\ell^{th}$ layer after applying the transformation function $f(\cdot)$. The output of the final layer, $h^{(L)}$, is typically used for downstream tasks, such as classification or natural language generation.
@@
1:Model $M$ with $L$ layers, Number of layers to prune $n$, Dataset $D$
2:Pruned model $M^{\prime}$
@@
To determine which layers to prune, we calculate the angular distance between the activation outputs of layer $\ell$ and layer $\ell+n$. For each potential starting layer $\ell$, the angular distance $d(h^{(\ell)}(D),h^{(\ell+n)}(D))$ is computed using a representative dataset $D$, which may be a representative pre-training dataset or one that is tailored to a specific downstream task. In our work, we use RedPajama (Computer, 2023) as the representative dataset to evaluate the sample distances. The angular distance metric is formally defined as,
$d(h^{(\ell)}(D),h^{(\ell+n)}(D))\equiv\frac{1}{\pi}\arccos\left(\frac{h^{(\ell)}_{T}(D)\cdot h^{(\ell+n)}_{T}(D)}{\left\lVert h^{(\ell)}_{T}(D)\right\rVert\left\lVert h^{(\ell+n)}_{T}(D)\right\rVert}\right),$ (1)
@@
where $\ell^{\star}$ corresponds to the layer with the smallest angular distance to its corresponding $n$-th successor layer. This optimization identifies a block of layers that exhibit high redundancy, as measured by their similar output activations. Pruning such a block is expected to have minimal impact on the model’s overall capacity. Once identified, layers from $\ell^{\star}$ to $\ell^{\star}+n-1$ are removed, and the model is restructured by directly connecting the output of layer $\ell^{\star}$ to the input of layer $\ell^{\star}+n$. This pruning operation can be efficiently implemented in deep learning frameworks such as PyTorch (Paszke et al., 2019), where the layers are typically encapsulated within a ModuleList or similar data structure. The layers within the identified block are excluded when defining the new, pruned model architecture.
Figure 2: Comparison of Llama3.1-8B Instruct using (left) angular cosine and (center) block influence (BI) score metrics. (right) Distribution of embedding similarities on GSM8k after fine-tuning on OpenMathInstruct for a pruned Llama3.1-8B Instruct model (i.e., 26 decoder layers, prune block size $n$ = 6). Self-data distilled fine-tuning (Self-Data FT) achieves higher similarity to the original baseline, indicating reduced distribution shift compared to standard supervised fine-tuning (SFT).
@@ 2.2 Self-Data Distillation for Pruned Models
After pruning a Transformer, fine-tuning is typically required to adapt the pruned model to specific downstream tasks. However, the fine-tuning process can amplify catastrophic forgetting, especially when the fine-tuning data distribution diverges from the original training distribution. To address this, we propose self-data distilled fine-tuning, which aligns the fine-tuning dataset with the original model’s learned distribution, helping to mitigate forgetting and maintain model quality across tasks.
@@
3 Ablation Studies on Layer-Pruning and Self-Data Distillation
@@
Effect of Layer Importance Metric
@@
We investigated two layer importance metrics, BI and angular cosine distance, to guide pruning decisions in the Llama3.1-8B Instruct model (see Figure 2). Both metrics assess the redundancy of layers by measuring the cosine distance between their inputs and outputs but differ in their focus. BI quantifies the change in hidden states across layers, while angular cosine distance emphasizes output similarity between consecutive layers. Despite minor differences in middle layers, both metrics produced comparable pruning results across block sizes. We ultimately chose the angular cosine metric due to its computational efficiency, leveraging cosine similarity without the added complexity of BI. This efficiency makes it more scalable for large models, and its direct measure of output similarity aligns with our intuition that layers producing highly similar outputs are likely redundant, making it a more practical tool for structured pruning in this study.
@@
Table 1: Model quality results for pruned Llama3.1-8B Instruct models across various pruning block sizes and fine-tuning strategies. Average accuracy is reported on the OpenLLM Leaderboard v1, along with the recovery percentage relative to the baseline model. Self-data distillation consistently outperforms other methods, particularly when combined with model merging (MM).
@@
Prune Block Size Model Savings Fine-tuning Method Dataset ARC-C (25-shot) HellaSwag (10-shot) TruthfulQA (0-shot) MMLU (5-shot) Winogrande (5-shot) GSM8k (5-shot) Avg. Score Avg. Recovery Baseline - No FT 60.92 80.16 54.02 68.15 77.58 75.59 69.40 - 4 10.86% (7.16B) No FT 55.20 75.70 52.40 67.79 75.29 56.18 64.06 92.31% SFT OpenMathInstruct 51.62 71.70 49.40 61.65 73.25 44.35 58.66 84.52% Self-Data Distillation OpenMathInstruct 53.93 74.27 50.61 65.34 74.90 69.44 64.75 93.29% Self-Data Distillation + MM OpenMathInstruct + Alpaca 56.22 76.48 51.60 66.93 74.43 69.37 65.84 94.86% 6 16.30% (6.72B) No FT 49.49 68.72 53.63 67.42 70.40 1.29 51.82 74.67% SFT OpenMathInstruct 46.93 68.51 50.81 59.98 70.01 43.82 56.68 81.66% Self-Data Distillation OpenMathInstruct 50.00 71.57 53.14 64.96 73.64 66.64 63.33 91.24% Self-Data Distillation + MM OpenMathInstruct + Alpaca 54.52 73.58 53.57 66.35 74.82 65.67 64.75 93.30% 8 21.73% (6.29B) No FT 44.71 61.22 56.11 65.57 65.98 0.00 48.93 70.50% SFT OpenMathInstruct 42.15 64.49 54.69 60.65 66.38 29.64 53.00 76.37% Self-Data Distillation OpenMathInstruct 46.67 65.70 53.32 64.87 71.27 57.70 59.92 86.38% Self-Data Distillation + MM OpenMathInstruct + Alpaca 46.93 67.85 53.33 65.81 72.30 56.41 61.24 88.24% 10 27.16% (5.85B) No FT 37.46 54.45 55.06 64.09 67.25 0.00 46.39 66.83% SFT OpenMathInstruct 39.33 57.38 51.47 57.44 64.48 15.39 47.58 68.56% Self-Data Distillation OpenMathInstruct 40.88 61.11 53.46 64.54 70.88 44.58 55.91 80.56% Self-Data Distillation + MM OpenMathInstruct + Alpaca 41.72 63.30 52.87 65.57 70.33 42.26 56.08 80.70%
Analysis on Self-Data Distilled Datasets

### v3

@@ Abstract
Large language models have driven significant progress in natural language processing, but their deployment requires substantial compute and memory resources. As models scale, compression techniques become essential for balancing model quality with computational efficiency. Structured pruning, which removes less critical components of the model, is a promising strategy for reducing complexity. However, one-shot pruning often results in significant quality degradation, particularly in tasks requiring multi-step reasoning. To recover lost quality, supervised fine-tuning (SFT) is commonly applied, but it can lead to catastrophic forgetting by shifting the model’s learned data distribution. Therefore, addressing the degradation from both pruning and SFT is essential to preserve the original model’s quality. In this work, we utilize self-data distilled fine-tuning to address these challenges. Our approach leverages the original, unpruned model to generate a distilled dataset that preserves semantic richness and mitigates catastrophic forgetting by maintaining alignment with the base model’s knowledge. Empirically, we demonstrate that self-data distillation consistently outperforms standard SFT, improving average accuracy by up to 8% on the HuggingFace OpenLLM Leaderboard v1. Specifically, when pruning six decoder blocks on Llama3.1-8B Instruct (i.e., 32 to 26 layers, reducing the model size from 8.03B to 6.72B parameters), our method retains 91.2% of the original model’s accuracy compared to 81.7% with SFT, while reducing real-world FLOPs by 16.3%. Furthermore, combining self-data distilled models through model merging yields enhanced quality retention. Additionally, leveraging these pruned models in speculative decoding increases token acceptance rates, thereby improving inference efficiency in applied settings.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. *Work done during an internship at Cerebras. Correspondence to: Anonymous Author <anon.email@domain.com>.
@@ 1 Introduction
The advent of large language models (LLMs) such as GPT-4 (OpenAI et al., 2024), Gemini (Gemini et al., 2024), and Llama 3 (Dubey et al., 2024) has revolutionized natural language processing (NLP), driving significant advancements across various tasks through extensive pre-training on textual data. These models, enhanced by supervised fine-tuning (SFT), demonstrate impressive instruction-following abilities (Ouyang et al., 2022; Touvron et al., 2023a), but come with high compute costs for both training and inference (Kaplan et al., 2020; Hoffmann et al., 2022). To address diverse deployment requirements across varying model scales, sizes, and compute budgets, compressing models for efficient inference is essential, particularly given the significant time, data, and resource constraints associated with training multiple multi-billion parameter models from scratch.
Most model compression techniques can be grouped into four main categories: knowledge distillation (KD) (Hinton et al., 2015), factorization (Hu et al., 2022), pruning (LeCun et al., 1989), and quantization (Han et al., 2015). In our work, we focus on pruning, though we aim for our method to inspire further developments across these other compression methods. Structured pruning, which selectively removes less critical components of a neural network, has emerged as a promising method for improving LLM efficiency (Ma et al., 2023). This method has gained attention for its ability to reduce memory and compute requirements, making inference more efficient. Recent works have shown that LLMs exhibit significant redundancy, particularly in the middle layers, where removing these layers has a minimal impact on overall model quality (Men et al., 2024; Gromov et al., 2024). The residual stream of the Transformer (Vaswani et al., 2017) architecture is only slightly modified by the output of non-essential layers, enabling the removal of these layers without drastically harming model quality.
@@
Despite its potential advantages, depth-wise structured pruning presents inherent challenges. It often leads to accuracy degradation, especially on tasks requiring multi-step reasoning, such as ARC-C (Clark et al., 2018) or GSM8k (Cobbe et al., 2021), where the structured order of layer outputs plays a crucial role. In these cases, pruning disrupts the flow of information between layers, resulting in poor model quality even after supervised fine-tuning (SFT) (Sun et al., 2024). While SFT can help recover some of the lost quality, it is generally insufficient for tasks with high reasoning complexity, where the structured sequence of layer outputs is essential. In addition, fine-tuning can amplify catastrophic forgetting (McCloskey & Cohen, 1989; Kotha et al., 2024), where the model loses previously learned information, particularly on tasks not represented in the fine-tuning data. Standard mitigation strategies, such as data replay (Ostapenko et al., 2022) or parameter importance-based methods (Kirkpatrick et al., 2017), often become impractical for LLMs due to their scale. Moreover, fine-tuning often leads to distribution shifts, further degrading model quality (Yang et al., 2024). As LLMs continue to grow in size and complexity, developing more effective strategies to mitigate these challenges during pruning is critical to unlocking its full potential.
@@
In our work, we propose a novel approach to mitigate the adverse effects of structured pruning by employing self-data distilled fine-tuning. Our method leverages the original, unpruned model as a seed language model to generate a distilled dataset that upholds semantic equivalence with the original task dataset. This approach not only preserves the semantic richness of the data but also mitigates catastrophic forgetting, a phenomenon where fine-tuned models lose their general instruction-following abilities due to the distribution shift introduced during standard SFT. As seen in Figure 1, we show that self-data distillation improves accuracy recovery by greater than 10% over SFT on the HuggingFace OpenLLM Leaderboard v1 (Beeching et al., 2023). Specifically, when pruning 6 blocks from Llama3.1-8B Instruct, our approach retains 91.2% of the original model’s accuracy compared to 81.7% with SFT, while also reducing FLOPs by 16.30%. In addition to improving the quality of pruned models, our work explores extending self-data distilled fine-tuning to speculative decoding (Leviathan et al., 2023; Chen et al., 2023). This approach allows pruned models to not only recover their accuracy but also achieve efficient inference through parallel token generation. The integration of self-data distillation with speculative decoding is a natural extension, as it leverages the semantic alignment achieved during data distillation to enhance the speculative capabilities of pruned models. By aligning the pruned model’s predictions more closely with the target model, speculative decoding can benefit from improved token acceptance rates, reducing latency while preserving model quality. The main contributions of our work are:
@@
  • To our knowledge, we are the first to introduce self-data distillation as a fine-tuning method for recovering the model quality of pruned models. Empirically, we show that self-data distillation on Llama3.1-8B Instruct and Mistral-7B-v0.3 Instruct (Jiang et al., 2023) consistently outperforms SFT across all pruned models.
@@
  • We extend self-data distilled fine-tuning to speculative decoding, showing that it enhances token acceptance rates and reduces latency during inference. This extension highlights the broader applicability of our approach, supporting both compression and acceleration of large language models.
@@ 2.1 Layer-Pruning Algorithm for Language Models
Transformers (Vaswani et al., 2017) have become a foundational architecture in deep learning, particularly for tasks involving natural language processing and sequence modeling. A standard Transformer consists of $L$ layers, each of which includes a multi-head self-attention mechanism followed by a feedforward network. These layers sequentially transform an input sequence into increasingly abstract representations, with each layer’s output serving as the input to the subsequent layer. Let $x^{(\ell)}$ denote the input to the $\ell^{th}$ layer, and $h^{(\ell)}=f(x^{(\ell)})$ denote the output of the $\ell^{th}$ layer after applying the transformation function $f(\cdot)$. The output of the final layer, $h^{(L)}$, is typically used for downstream tasks (e.g., classification or natural language generation).
@@
1:  Input: Model $M$ with $L$ layers, Number of layers to prune $n$, Dataset $D$
2:  Output: Pruned model $M^{\prime}$
@@
To determine which layers to prune, we calculate the angular distance between the activation outputs of layer $\ell$ and layer $\ell+n$. For each potential starting layer $\ell$, the angular distance $d(h^{(\ell)}(D),h^{(\ell+n)}(D))$ is computed using a representative dataset $D$, which may be a representative pre-training dataset or one that is tailored to a specific downstream task. In our work, we use RedPajama (Computer, 2023) as the representative dataset to evaluate the sample distances (see more details in Appendix A.3). The angular distance metric, $d(h^{(\ell)}(D),h^{(\ell+n)}(D))$, is formally defined as,
$\frac{1}{\pi}\arccos\left(\frac{h^{(\ell)}_{T}(D)\cdot h^{(\ell+n)}_{T}(D)}{\left\lVert h^{(\ell)}_{T}(D)\right\rVert\left\lVert h^{(\ell+n)}_{T}(D)\right\rVert}\right),$ (1)
@@
where $\ell^{\star}$ corresponds to the layer with the smallest angular distance to its corresponding $n$-th successor layer. This optimization identifies a block of layers that exhibit high redundancy, as measured by their similar output activations. Pruning such a block is expected to have minimal impact on the model’s overall capacity. Once identified, layers from $\ell^{\star}$ to $\ell^{\star}+n-1$ are removed, and the model is restructured by directly connecting the output of layer $\ell^{\star}$ to the input of layer $\ell^{\star}+n$.
@@ 2.2 Self-Data Distillation for Pruned Models
After pruning a Transformer, fine-tuning is typically required to adapt the pruned model to specific downstream tasks. However, the fine-tuning process can amplify catastrophic forgetting, especially when the fine-tuning data distribution diverges from the original training distribution. To address this, we utilize self-data distilled fine-tuning (Yang et al., 2024), which aligns the fine-tuning dataset with the original model’s learned distribution, to mitigate forgetting and maintain model quality across a diverse range of tasks.
@@
Figure 3: Comparison of Llama3.1-8B Instruct using (left) angular cosine and (right) block influence (BI) score metrics. Both metrics highlight inherent redundancy in the middle layers of Llama3.1-8B Instruct, suggesting that these layers can be pruned with minimal impact on overall model quality.
Figure 4: Distribution of embedding similarities on GSM8k test dataset. We present the distribution of embedding similarities after fine-tuning a structurally pruned variant of the Llama3.1-8B Instruct model on 50k samples of OpenMathInstruct. We compute the cosine similarity between the sentence embeddings of the pruned models and those generated by the original Llama3.1-8B Instruct model. The plots show: (left) 28 decoder layers (prune block size = 4), (center) 26 decoder layers (prune block size = 6), and (right) 24 decoder layers (prune block size = 8). Self-Data Distilled Fine-Tuning (Self-Data FT) achieves higher similarity to the original baseline model, indicating a reduced distribution shift compared to Supervised Fine-Tuning (SFT).
3 Ablation Studies
@@
3.1 Effect of Layer Importance Metric
@@
We investigated two layer importance metrics, BI and angular cosine distance, to guide pruning decisions in the Llama3.1-8B Instruct model (see Figure 4). Both metrics assess the redundancy of layers by measuring the cosine distance between their inputs and outputs but differ in their focus. BI quantifies the change in hidden states across layers, while angular cosine distance emphasizes output similarity between consecutive layers. Despite minor differences in the middle layers, both metrics produced comparable pruning results across block sizes. We ultimately chose the angular cosine metric for its computational efficiency, as it uses cosine similarity on only the last token, whereas BI operates over the entire sequence, adding computational overhead. This efficiency makes it more scalable for LLMs, and its direct measure of output similarity aligns with our intuition that layers producing highly similar outputs are likely redundant, making it a practical tool for structured pruning in this study.
@@
Table 1: Model quality results for pruned Llama3.1-8B Instruct models across various pruning block sizes and fine-tuning strategies. Average accuracy is reported on the OpenLLM Leaderboard v1, along with the recovery percentage relative to the baseline model. Self-data distillation consistently outperforms outperforms standard supervised fine-tuning (SFT).
@@
Prune Block Size Model Savings Fine-tuning Method Dataset ARC-C (25-shot) HellaSwag (10-shot) TruthfulQA (0-shot) MMLU (5-shot) Winogrande (5-shot) GSM8k (5-shot) Avg. Score Avg. Recovery Baseline - No FT 60.92 80.16 54.02 68.15 77.58 75.59 69.40 - 2 5.23% (7.61B) No FT 58.71 78.12 53.21 63.15 76.72 67.39 66.22 95.41% SFT OpenMathInstruct 52.91 73.24 51.77 60.91 73.75 44.95 59.59 85.85% Self-Data Distillation OpenMathInstruct 56.41 75.03 50.23 66.17 75.18 69.97 65.51 94.37% 4 10.86% (7.16B) No FT 55.20 75.70 52.40 67.79 75.29 56.18 64.06 92.31% SFT OpenMathInstruct 51.62 71.70 49.40 61.65 73.25 44.35 58.66 84.52% Self-Data Distillation OpenMathInstruct 53.93 74.27 50.61 65.34 74.90 69.44 64.75 93.29% 6 16.30% (6.72B) No FT 49.49 68.72 53.63 67.42 70.40 1.29 51.82 74.67% SFT OpenMathInstruct 46.93 68.51 50.81 59.98 70.01 43.82 56.68 81.66% Self-Data Distillation OpenMathInstruct 50.02 71.57 53.14 64.96 73.64 66.64 63.33 91.24% 8 21.73% (6.29B) No FT 44.71 61.22 56.11 65.57 65.98 0.00 48.93 70.50% SFT OpenMathInstruct 42.15 64.49 54.69 60.65 66.38 29.64 53.00 76.37% Self-Data Distillation OpenMathInstruct 46.67 65.70 53.32 64.87 71.27 57.70 59.92 86.38% 10 27.16% (5.85B) No FT 37.46 54.45 55.06 64.09 67.25 0.00 46.39 66.83% SFT OpenMathInstruct 39.33 57.38 51.47 57.44 64.48 15.39 47.58 68.56% Self-Data Distillation OpenMathInstruct 40.88 61.11 53.46 64.54 70.88 44.58 55.91 80.56%
Table 2: Model quality results for pruned Mistral-7B-v0.3 Instruct models across various pruning block sizes and fine-tuning strategies. Average accuracy is reported on the OpenLLM Leaderboard v1, along with the recovery percentage relative to the baseline model. Self-data distillation consistently outperforms standard supervised fine-tuning (SFT).
Prune Block Size Model Savings Fine-tuning Method Dataset ARC-C (25-shot) HellaSwag (10-shot) TruthfulQA (0-shot) MMLU (5-shot) Winogrande (5-shot) GSM8k (5-shot) Avg. Score Avg. Recovery Baseline - No FT 63.39 84.52 59.65 61.99 78.37 48.75 66.11 - 2 6.02% (6.83B) No FT 60.24 62.31 58.57 62.11 78.77 32.75 59.12 89.43% SFT OpenMathInstruct 54.01 79.38 53.71 57.61 74.82 34.04 58.93 89.13% Self-Data Distillation OpenMathInstruct 53.93 74.27 50.61 65.34 74.90 69.44 64.75 93.29% 4 12.03% (6.39B) No FT 57.17 78.89 57.44 61.36 77.19 9.33 56.90 86.06% SFT OpenMathInstruct 52.65 75.29 53.16 58.21 74.51 27.22 56.84 85.97% Self-Data Distillation OpenMathInstruct 55.29 75.77 53.33 60.12 75.22 34.65 59.06 89.34% 6 18.06% (5.96B) No FT 50.25 72.67 55.85 61.56 75.45 0.83 52.77 79.82% SFT OpenMathInstruct 47.52 70.15 51.86 58.29 74.91 20.47 53.87 81.47% Self-Data Distillation OpenMathInstruct 48.98 70.67 52.28 59.56 72.22 28.13 55.31 83.65% 8 24.08% (5.52B) No FT 41.04 65.27 58.10 61.05 71.74 0.00 49.53 74.92% SFT OpenMathInstruct 43.09 66.11 52.74 58.52 72.21 9.48 50.35 76.16% Self-Data Distillation OpenMathInstruct 45.14 67.09 53.61 59.81 71.43 19.03 52.68 79.69% 10 30.10% (5.08B) No FT 34.30 52.14 58.67 35.76 65.11 0.00 41.00 62.01% SFT OpenMathInstruct 39.33 59.07 52.07 57.49 68.75 6.07 47.13 71.29% Self-Data Distillation OpenMathInstruct 39.76 59.25 51.61 58.42 69.61 14.39 48.84 73.87%
3.2 Analysis on Self-Data Distilled Datasets